Royalshani222:

Find the length of the base of a square pyramid if the volume is 576 cubic inches and has a height of 3 inches.

kittybasil:

Volume of a square-base pyramid: \(\frac{1}{3}\cdot bh\) where \(b\) is the area of the Base (\(l\times w\)) or length times width; and \(h\) is the height (3 inches, as given).

kittybasil:

**Forgot to put the volume in the original equation. So the volume formula is actually \(V=\frac{1}{3}bh\) Moving on... As the base is a square, it is equilateral and all lengths/widths are the same. Therefore, we can also rewrite the formula with length as variable \(x\): where \(b=l\times w\to x\cdot x\to x^{2}\)\[V=\frac{1}{3}\cdot\color{red}{b}h\]thus becomes \(V=\frac{1}{3}\cdot\color{red}{x^{2}}h\), and when we input the \(\color{fuchsia}{\text{volume}}\) and \(\color{lime}{\text{height}}\) value we have\[\color{fuchsia}{576}=\frac{1}{3}\cdot x^{2}(\color{lime}{3})\]

kittybasil:

Now we solve for variable \(x\). Let's first see what we can cancel out. We have a 3 in the numerator (anything without a fraction form is considered numerator) and a 3 in the denominator (the bottom part of the fraction). These cancel out:\[576=\frac{1}{\cancel{3}}\cdot x^{2}(\cancel{3})\]\[576=x^{2}\]

kittybasil:

Now we need to isolate the variable to find its value. To do this, we will remove the square by square root.\[\sqrt{576}=\sqrt{x^{2}}\]\[24=x\]So your pyramid base has a length/width of 24 inches. Why inches? Because volume is in cubic units, and the value given for your volume is in cubic \(\text{inches}\).